5
E i(i-5)(i-3)
i=0
Find the sum of the series. I am not sure the steps to solve.

Answer by stanbon(75887)

5
E i(i-5)(i-3)
i=0
==================
If i = 0 the product is 0
If i = 1 the product is 1*-4*-2 = 8
If i = 2 the product is 2*-3*-1 = 6
If i = 3 the product is 3*-2*0 = 0
If i = 4 the product is 4*-1*1 = -4
If i = 5 the product is 5*0*2 = 0
----
The sum = 8+6+0-4+0 = 10
=====================================
Cheers,
Stan H.
========